20. In year 0, Country A has a real GDP per capita of $1,200. If Country A grows at a constant rate of 2% per year and Country A's population remains constant, what is Country A's real GDP per capita by year 20? (Round to the nearest dollar.)
Answer #1

ANSWER:

Pv = $1,200

n = 20 years

i = 2%

fv = pv * (1 + r) ^ n

fv = 1,200 * (1 + 2%) ^ 20

fv = 1,200 * (1.02) ^ 20

fv = 1,200 * 1.49

fv = 1,783.14

so the gdp per capita after 20 years will be $1,783.14